Find three consecutive odd integers such that the sum of the first and second is 27 less than 3 times the third
Answer by Alan3354(69443)

Find three consecutive odd integers such that the sum of the first and second is 27 less than 3 times the third
======================
Similar problem.
---------
find three consecutive positive odd integers such that the product of the two smaller exceeds the largest by 26
-------------
Use n-1, n+1 and n+3
n^2-1 = n+3 + 26
n^2 - n - 30 = 0
etc